# Foundational theories of hesitant fuzzy sets and hesitant fuzzy information systems and their applications for multi-strength intelligent classifiers

### v3

@@ Abstract
Hesitant fuzzy sets are widely used in certain instances of uncertainty and hesitation. In sets, the inclusion relationship is an important and foundational definition. Thus, as a kind of set, hesitant fuzzy sets require an explicit definition of inclusion relationship. Based on the hesitant fuzzy membership degree of discrete form, several kinds of inclusion relationships for hesitant fuzzy sets are proposed in this work. Then, some foundational propositions of hesitant fuzzy sets are presented, along with propositions of families of hesitant fuzzy sets. Some foundational propositions of hesitant fuzzy information systems are proposed with respect to parameter reductions and an example and an algorithm are given to illustrate the processes of parameter reduction. Finally, a multi-strength intelligent classifier is proposed to make health state diagnoses for complex systems.
@@
Hesitant fuzzy sets, hesitant fuzzy information systems, parameter reductions, multi-strength intelligent classifier
@@ 1 Introduction
Set theory established by Cantor FP (hereafter referred to as classical set) is the foundation of modern mathematics. Meanwhile, modern mathematics, such as group theory EJB ; DDJ , topology SWA , graph theory GHH and so on, are the foundations of computer science. After classical set theory was perfected, other kinds of sets were developed, such as fuzzy set ZADE , interval-valued fuzzy set GOMB , intuitionistic fuzzy set AKT and hesitant fuzzy set TV and so on. Several kinds of sets had developed their foundational theoretical systems except hesitant fuzzy set. Especially, one of the most foundational definitions, the definition of the inclusion relationship between two hesitant fuzzy sets, is not sufficiently clear. This article develops the foundational theories of hesitant fuzzy sets and hesitant fuzzy information systems, and proposes a multi-strength intelligent classifier to make health state diagnoses for complex systems.
For the classical set FP , a set describes the membership degree of an element in two cases, namely, 0% and 100%. The relationship between an element and a set is determined depending on the element in or not in the set. However, there are uncertainties in the real world. To deal with the uncertain cases, Zadeh proposed the concept of fuzzy set ZADE wherein the membership degree of an element of a set is a value in $[0,1]$. The membership degree of a fuzzy set represents the probability that an element belongs to a set TZH ; YSB and is no longer just 0% and 100%. Subsequently, Gorzalczany GOMB proposed the concept of interval-valued fuzzy set in which the membership degree of an element of a set is an interval included in $[0,1]$. Interval-valued fuzzy set can deal with cases wherein we cannot determine the exact probability of an element belonging to a set, but we can determine an interval that the probability must appear in this interval BSO ; FKM . Furthermore, Atanassov AKT proposed the concept of intuitionistic fuzzy set wherein the membership degree of an element of a set is a bibasic array. Intuitionistic fuzzy set can describe things with a dual character, one number of the bibasic array expresses the positive aspect and the other number expresses the negative aspect KRA ; PLDY .
Torra TV proposed the concept of hesitant fuzzy set wherein the membership degree of an element of a set is a multidimensional array. Hesitant fuzzy sets focus on cases of uncertainty and hesitation, which are common in the real world. For instance, about the data sampling for equipment monitoring and decision scoring by an expert team, the feedback of such cases always comes in the form of arrays, and the multiple different values can make us indecisive. Thus, hesitant fuzzy sets with a membership degree of a discrete array are more suitable in describing the indecisiveness in decisions compared with the classical, fuzzy, interval-valued fuzzy, and intuitionistic fuzzy sets. Hesitant fuzzy sets are widely used in different fields, such as decision-making HJH ; LLYX ; LNZQY ; GXLY , risk evaluation HXZS , agglomerative hierarchical clustering JLGL , investment policy DYM , modeling for complex product manufacturing tasks BLYY , linguistic perceptual AWNJE , forecasting HLJW ; PMNME , and so on.
@@
However, there is no sufficiently clear definition of the inclusion relationship between two hesitant fuzzy sets. The definition of inclusion relationship is a significant foundation for sets, because if a set is a subset of another set and vice versa, then these two sets are equivalent. Based on the hesitant fuzzy membership degree of discrete form, we propose in this work several kinds of inclusion relationships for hesitant fuzzy sets and then present some foundational propositions of hesitant fuzzy sets and the families of hesitant fuzzy sets. It is worth noting that some foundational propositions of classical sets do not hold in hesitant fuzzy sets.
In recent years, information systems with uncertain information have become a hot topic due to the rising instances of uncertain information in the real world. Plenty of studies have been conducted on this topic, such as fuzzy information system NXSZ , fuzzy incomplete information system WWZJ , interval-valued information system DJWZ ; SDZX , and intuitionistic fuzzy information system HBLH . Based on the foundational propositions of hesitant fuzzy sets, we then study some theories with respect to parameter reductions in hesitant fuzzy information systems.
Analyzing high-dimensional data with small sample sizes is a challenge, the health status of a complex system with such data is difficult to diagnose. Based on the foundational propositions of hesitant fuzzy information systems, a multi-strength intelligent classifier in knowledge bases PALA ; PASA is proposed to make health state diagnoses for complex systems. This multi-strength intelligent classifier is a candidate tool for dimensionality reduction of high-dimensional data, classification, diagnosis, evaluation, decision-making, and so on.
The remainder of the paper is constructed as follows. Section 2 presents several kinds of inclusion relationships of hesitant fuzzy sets, while some foundational propositions of hesitant fuzzy sets are discussed, showing some rules that hold in classical sets or hesitating fuzzy sets. Section 3 presents some propositions on families of hesitant fuzzy sets. Section 4 proposes some propositions with respect to parameter reductions in hesitant fuzzy information systems and provides an algorithm for parameter reductions. Section 5 proposes a multi-strength intelligent classifier in knowledge bases to make health state diagnoses for complex systems. Section 6 presents a conclusion.
@@ 2 Foundations of hesitant fuzzy sets
In this section, we briefly recall some basic concepts about hesitant fuzzy sets and then present several kinds of inclusion relationships of hesitant fuzzy sets that are suitable for hesitant fuzzy membership degree of discrete form. We focus on whether some rules of classical sets hold in hesitant fuzzy sets. In subsequent studies of hesitant fuzzy sets, researchers must avoid the intuitions of classical sets as well as the rules held in the classical sets but not in the hesitant fuzzy sets.
If two sets $A$ and $B$ are classical sets, $A\sqcap B$ and $A\sqcup B$ represent the intersection and union of $A$ and $B$, respectively. Furthermore, $A\sqsubset B$ represents that $A$ is a subset of the classical set $B$. Let $U$ be a universal set and $E$ be a set of parameters. Let $HF(U)$ be the set of all hesitant fuzzy sets defined over $U$.
@@ 2.1 Reviews of fuzzy sets and hesitant fuzzy sets
ZADE A fuzzy set $F$ on $U$ is a mapping $F:U\rightarrow[0,1]$.
@@
ZADE $F_{1}$ and $F_{2}$ are two fuzzy sets on $U$, $F_{1}$ is a subset of $F_{2}$ if $F_{1}(x)\leqslant F_{2}(x)$ for all $x\in U$, denoted as $F_{1}\subset F_{2}$.
@@
TV A hesitant fuzzy set on $U$ is defined as a function that when applied to $U$ returns a subset of $[0,1]$.
@@
TV For each $x\in U$, and a hesitant fuzzy set $H$, the lower bound and upper bound of $H(x)$ are defined
lower bound $H^{-}(x)=min\ H(x)$,
upper bound $H^{+}(x)=max\ H(x)$.
@@
TV Given two hesitant fuzzy sets represented by their membership functions $H_{1}$ and $H_{2}$, their union and intersection are defined
union $(H_{1}\cup H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\geqslant max(H_{1}^{-}(x),H_{2}^{-}(x))\}$,
intersection $(H_{1}\cap H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\leqslant min(H_{1}^{+}(x),H_{2}^{+}(x))\}$.
Let $U=\{x,y\}$, there are three common expression forms of hesitant fuzzy set, such as $A=\{\frac{0.6,0.5,0.3}{x},\frac{0.5,0.3,0.2}{y}\}$, $A=\frac{0.6,0.5,0.3}{x}+\frac{0.5,0.3,0.2}{y}$ and $A=\{<x,(0.6,0.5,0.3)>,<y,(0.5,0.3,0.2)>\}$. $A=B$ means that $A$ and $B$ are perfectly consistent in this article, i.e., $B=\frac{0.3,0.6,0.5}{x}+\frac{0.2,0.5,0.3}{y}$.
@@
2.2 Inclusion relationships of hesitant fuzzy sets
@@
For two hesitant fuzzy sets $H_{1}$ and $H_{2}$, Babitha et al BJ introduced that $H_{1}$ is a hesitant fuzzy subset of $H_{2}$ if $H_{1}(x)\subset H_{2}(x)$ for all $x\in U$. However, what is the meaning of $H_{1}(x)\subset H_{2}(x)$? There is no detailed description in BJ .
Carlos et al JCVT introduced a definition of inclusion relationship for hesitant fuzzy sets, i.e., $H_{1}\subset H_{2}$Leftrightarrow $H_{1}\cup H_{2}=H_{2}$. By the description in JCVT , $H_{1}(x)\subset H_{2}(x)\Leftrightarrow\{h:h\in H_{2}(x)\}=H_{2}(x)=(H_{1}\cup H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\geqslant\max(H_{1}^{-}(x),H_{2}^{-}(x))\}$, i.e., $H_{1}(x)\subset H_{2}(x)$ implies $h<H_{2}^{-}(x)$ for all $h\in H_{1}(x)$. The definition of inclusion relationship for hesitant fuzzy sets in JCVT is not perfect enough for these aspects, shown as (i) and (ii),
@@
(i) The inclusion relationship of sets is also the ordering relationship of sets. A foundational and important proposition of inclusion relationship, $A\subset B$ and $B\subset A$ if and only if $A=B$, i.e., the equivalent sets contain each other. However, by the definition of inclusion relationship in JCVT , if $H_{1}(x)\subset H_{2}(x)$ and $H_{2}(x)\subset H_{1}(x)$, then a contradiction is produced, i.e., $H_{1}^{+}(x)<H_{2}^{-}(x)<H_{2}^{+}(x)<H_{1}^{-}(x)<H_{1}^{+}(x)$.
@@
(ii) The definition of inclusion relationship in JCVT is too specific to have a small range of applications. For example, let $H_{1}(x)=\{0.1,0.2,0.5\}$, $H_{2}(x)=\{0.6,0.7,0.9\}$ and $H_{3}(x)=\{0.1,0.5,0.8\}$, then $H_{1}(x)\subset H_{2}(x)$ is obvious. However, this definition cannot describe the relationship of $H_{2}(x)$ and $H_{3}(x)$ and is not applicable for many cases of hesitant fuzzy sets.
In this work, we propose several kinds of inclusion relationships for hesitant fuzzy sets based on the strength of the information in hesitant fuzzy membership degrees. First, we present a detailed example to illustrate the Definition 2.7.
@@
Let $U=\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}\}$ be a set of decision-making schemes and $H$ be an expert team that consists of three experts. $H(U)=\frac{0.9,0.2}{x_{1}}+\frac{0.6,0.6,0.5}{x_{2}}+\frac{0.7,0.5,0.5}{x_{3}}+\frac{0.8,0.6,0.5}{x_{4}}+\frac{0.9,0.3,0.1}{x_{5}}+\frac{0.9,0.8,0.7}{x_{6}}$ are the schemes’ scores that are given by experts, in which the scores for $x_{1}$ are 0.9 and 0.2 that are obtained through the evaluations made by experts. One of the three experts fails to evaluate the scheme $x_{1}$.
(1) Here, $x_{1}$ has a score of 0.9, which is greater than or equal to all the scores for $x_{2}$, it is possible that the scheme $x_{1}$ is better than the scheme $x_{2}$, denoted as $H(x_{2})\subset_{p}H(x_{1})$.
(2) Here, $0.5670.55=mean[H(x_{1})]<mean[H(x_{2})]=$ 0.567, where $mean[\cdot]$ is the mean value operator. In comparing the mean values of scores, we find that the scheme $x_{2}$ is better than the scheme $x_{1}$, denoted as $H(x_{1})\subset_{m}H(x_{2})$.
(3) On the one hand, the best score of $x_{3}$ is 0.7, which is greater than or equal to the best score of $x_{2}$. On the other hand, the worst score of $x_{3}$ is 0.5, which is greater than or equal to the worst score of $x_{2}$. To compare the respective best and worst cases of schemes $x_{2}$ and $x_{3}$, it is acceptable that the scheme $x_{3}$ is better than the scheme $x_{2}$, denoted as $H(x_{2})\subset_{a}H(x_{3})$.
(4) To compare the scores for schemes $x_{3}$ and $x_{4}$ one by one ($0.7\leqslant 0.8$; $0.5\leqslant 0.6$; and $0.5\leqslant 0.5$), it is strongly credible that the scheme $x_{4}$ is better than the scheme $x_{3}$, denoted as $H(x_{3})\subset_{s}H(x_{4})$.
(5) We can obtain $H(x_{1})\subset_{s}H(x_{5})$ after truncating the tail score of $x_{5}$, i.e., deleting the score 0.1 of $x_{5}$, this case is denoted as $H(x_{1})\subset_{st}H(x_{5})$ and is recorded briefly as $H(x_{1})\subset_{t}H(x_{5})$.
(6) The worst score of $x_{6}$ is greater than or equal to the best score of $x_{3}$. Thus, it is necessary that the scheme $x_{6}$ is better than the scheme $x_{3}$, which is denoted as $H(x_{3})\subset_{n}H(x_{6})$.
Comparing the mean value of scores is a common thought for decision-making; however, while doing so, some important information may be lost, such as the best and the worst scores.
@@
(2) Let $q$ be a positive integer. To take $q$ larger numbers of $W$ and construct a best $q$-subsequence of $W$, denoted as $W(q)$, i.e., $|W(q)|=q$ and $w\leqslant min(W(q))$ for $w\in W-W(q)$.
@@
(3) Let $q=min\{|H_{1}(x)|,|H_{2}(x)|\}$, $H_{1}(x)(q)$ and $H_{2}(x)(q)$ be the best $q$-subsequences of $H_{1}(x)$ and $H_{2}(x)$, respectively. Thereafter, $H_{1}(x)(q)\preccurlyeq H_{2}(x)(q)$ if and only if one of $H_{1}(x)\subset_{s}H_{2}(x)$ and $H_{1}(x)\subset_{t}H_{2}(x)$ holds.
(4) Let $K=H_{1}(x)\sqcap H_{2}(x)$, if $max(H_{1}(x)-K)\leqslant min(K)\leqslant max(K)\leqslant min(H_{2}(x)-K)$, then one of $H_{1}(x)\subset_{s}H_{2}(x)$ and $H_{1}(x)\subset_{t}H_{2}(x)$ holds.

### v4

# Foundational theories of hesitant fuzzy sets and families of hesitant fuzzy sets

@@ Abstract
Hesitant fuzzy sets find extensive application in specific scenarios involving uncertainty and hesitation. In the context of set theory, the concept of inclusion relationship holds significant importance as a fundamental definition. Consequently, as a type of sets, hesitant fuzzy sets necessitate a clear and explicit definition of the inclusion relationship. Based on the discrete form of hesitant fuzzy membership degrees, this study proposes multiple types of inclusion relationships for hesitant fuzzy sets. Subsequently, this paper introduces foundational propositions related to hesitant fuzzy sets, as well as propositions concerning families of hesitant fuzzy sets.
@@
Hesitant fuzzy sets , inclusion relationship , families of hesitant fuzzy sets
@@ 1 Introduction
The foundation of modern mathematics lies in set theory established by Cantor [1] (referred to as classical sets hereafter). Moreover, modern branches of mathematics, including group theory [2, 3], topology [4], graph theory [5], among others, form the underpinnings of computer science. Subsequent to the refinement of classical set theory, various types of sets emerged, encompassing fuzzy sets [6, 7], interval-valued fuzzy sets [8], intuitionistic fuzzy sets [9], and hesitant fuzzy sets [10], among others. With the exception of hesitant fuzzy sets, several types of sets have established their foundational theoretical frameworks. Notably, the definition of the inclusion relationship between two hesitant fuzzy sets, which is one of the fundamental definitions, lacks adequate clarity. This article develops the foundational theories of hesitant fuzzy sets and hesitant fuzzy information systems, and introduces a multi-strength intelligent classifier for diagnosing the health states of complex systems.
In the case of the classical set [1], the membership degree of an element is described as either 0% or 100%, and the element’s relationship with the set is determined by whether it belongs to the set or not. Nevertheless, the real world is fraught with uncertainties. To address such uncertain scenarios, Zadeh introduced the concept of fuzzy sets [6], where the membership degree of an element in a set is a value ranging from 0 to 1. The membership degree of a fuzzy set denotes the probability of an element belonging to the set [11], thereby extending beyond the binary values of 0% and 100%. Later, Gorzalczany [8] introduced the notion of interval-valued fuzzy sets, where the membership degree of an element in a set is represented by an interval included in $[0,1]$. Interval-valued fuzzy sets can handle situations where the precise probability of an element belonging to a set cannot be determined, but an interval can be identified within which the probability must lie [12]. Additionally, Atanassov [9] introduced the notion of intuitionistic fuzzy sets, where the membership degree of an element in a set is represented by a bipartite array. Intuitionistic fuzzy sets can capture entities with dual characteristics, with one value in the bipartite array representing the positive aspect and the other value representing the negative aspect [13, 14].
Torra [10] introduced the concept of hesitant fuzzy sets, where the membership degree of an element in a set is represented by a multidimensional array. Hesitant fuzzy sets are designed to address situations of uncertainty and hesitation, which are prevalent in real-world scenarios. For example, in scenarios involving data sampling for equipment monitoring and decision scoring by expert teams, the feedback often manifests as arrays with multiple diverse values, leading to indecisiveness. Consequently, hesitant fuzzy sets with a membership degree represented by a discrete array are better suited for capturing the indecisiveness in decision-making compared to classical, fuzzy, interval-valued fuzzy, and intuitionistic fuzzy sets. Hesitant fuzzy sets find extensive applications in various domains, including decision-making [15, 16], attribute reduction [17], classification [18], linguistic perceptual [19], and forecasting [20], among others.
@@
Nevertheless, a sufficiently clear definition of the inclusion relationship between two hesitant fuzzy sets is currently lacking. The definition of the inclusion relationship serves as a crucial foundation for sets, as it establishes equivalence between two sets when one is a subset of the other and vice versa. Based on the discrete form of hesitant fuzzy membership degrees, this study introduces multiple types of inclusion relationships for hesitant fuzzy sets and subsequently presents foundational propositions concerning hesitant fuzzy sets and their families. It is important to highlight that some foundational propositions applicable to classical sets do not hold in the case of hesitant fuzzy sets.
The remainder of this paper is organized as follows. Section 2 introduces various types of inclusion relationships among hesitant fuzzy sets, along with their foundational propositions, and shows whether certain rules that apply to classical sets also hold for hesitant fuzzy sets. In Section 3, we put forth propositions concerning families of hesitant fuzzy sets. Section 4 provides a conclusion.
@@ 2 Foundations of hesitant fuzzy sets
In this section, we provide a brief overview of fundamental concepts related to hesitant fuzzy sets and introduce various types of inclusion relationships that are applicable to the discrete form of hesitant fuzzy membership degrees. Our focus is on examining whether certain rules that apply to classical sets also hold for hesitant fuzzy sets. In further research on hesitant fuzzy sets, it is crucial for researchers to refrain from relying on the intuitions derived from classical sets and to be mindful of the rules that are valid in classical sets but not in hesitant fuzzy sets.
If two sets $A$ and $B$ are classical sets, $A\sqcap B$ and $A\sqcup B$ represent the intersection and union of $A$ and $B$, respectively. Furthermore, $A\sqsubset B$ represents that $A$ is a subset of the classical set $B$. Let $U$ be a universal set and $E$ be a set of parameters.
@@ 2.1 Reviews of fuzzy sets and hesitant fuzzy sets
[6] A fuzzy set $F$ on $U$ is a mapping $F:U\rightarrow[0,1]$.
@@
[6] $F_{1}$ and $F_{2}$ are two fuzzy sets on $U$, $F_{1}$ is a subset of $F_{2}$ if $F_{1}(x)\leqslant F_{2}(x)$ for all $x\in U$, denoted as $F_{1}\subset F_{2}$.
@@
[21] A hesitant fuzzy element is a non-empty, finite subset of $[0,1]$.
@@
[10] A hesitant fuzzy set on $U$ is defined as a function that when applied to $U$ returns a subset of $[0,1]$.
In the following, $HF(U)$ denotes the set of all hesitant fuzzy sets defined over $U$.
@@
[10] For each $x\in U$, and a hesitant fuzzy set $H$, the lower bound and upper bound of $H(x)$ are defined
lower bound $H^{-}(x)=inf\ H(x)$,
upper bound $H^{+}(x)=sup\ H(x)$.
Definition 2.6.
@@
[10] Given two hesitant fuzzy sets represented by their membership functions $H_{1}$ and $H_{2}$, their union and intersection are defined
union $(H_{1}\cup H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\geqslant sup(H_{1}^{-}(x),H_{2}^{-}(x))\}$,
intersection $(H_{1}\cap H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\leqslant inf(H_{1}^{+}(x),H_{2}^{+}(x))\}$.
Definition 2.7.
[10] For $x\in U$ and $H\in HF(U)$, the complement of $H$ is denoted as $H^{c}$, where
$H^{c}(x)=\sqcup_{\gamma\in H(x)}\{1-\gamma\}$.
Theorem 2.8.
[10, 21] The following statements hold for $A,B,C\in HF(U)$,
(1) $(A^{c})^{c}=A$.
(2) $(A\cap B)^{c}=A^{c}\cup B^{c}$.
(3) $(A\cup B)^{c}=A^{c}\cap B^{c}$.
(4) $A\cap B=B\cap A$, $A\cup B=B\cup A$.
(5) $(A\cap B)\cap C=A\cap(B\cap C)$, $(A\cup B)\cup C=A\cup(B\cup C)$.
Let $U=\{x,y\}$, $A(x)=\{0.6,0.5,0.3\}$ and $A(y)=\{0.5,0.3,0.2\}$. There are three common expression types for the hesitant fuzzy set $A$, shown as follows,
[…]
Furthermore, $A=B$ means that $A(x)$ and $B(x)$ are perfectly consistent for each $x\in U$. For example, $B=\frac{\{0.3,0.6,0.5\}}{x}+\frac{\{0.2,0.5,0.3\}}{y}$, then $B=A$; $C=\frac{\{0.3,0.3,0.6,0.5\}}{x}+\frac{\{0.2,0.5,0.3\}}{y}$, then $C\neq A$.
@@
2.2 The proposed inclusion definitions of hesitant fuzzy sets and their relationships
@@
For two hesitant fuzzy sets $H_{1}$ and $H_{2}$, Babitha et al [22] defined that $H_{1}$ is a hesitant fuzzy subset of $H_{2}$ if $H_{1}(x)\subset H_{2}(x)$ for all $x\in U$. However, the reference [22] lacks a detailed description of $H_{1}(x)\subset H_{2}(x)$.
For two fuzzy sets $F_{1}$ and $F_{2}$, we have $F_{1}\subset F_{2}\Leftrightarrow F_{1}\cup F_{2}=F_{2}$. Carlos et al [23] introduced a definition of inclusion relationship for hesitant fuzzy sets, i.e., $H_{1}\subset H_{2}$Leftrightarrow $H_{1}\cup H_{2}=H_{2}$. This definition (in [23]) inherits the idea of fuzzy sets, which is convenient for scholars to refer to the existing theory of fuzzy sets to study hesitant fuzzy sets. However, we hold two different views on this definition (in [23]), shown as (i) and (ii),
@@
(i) The inclusion relationship of sets is also the ordering relationship of sets. A foundational and important proposition of inclusion relationship, $A\subset B$ and $B\subset A$ if and only if $A=B$, i.e., the equivalent sets contain each other. We assume that $H_{1}=H_{2}$ means that $H_{1}(x)$ and $H_{2}(x)$ are perfectly consistent for each $x\in U$. By the description in [23], $H_{1}(x)\subset H_{2}(x)\Leftrightarrow\{h:h\in H_{2}(x)\}=H_{2}(x)=(H_{1}\cup H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\geqslant sup(H_{1}^{-}(x),H_{2}^{-}(x))\}$, i.e., $H_{1}(x)\subset H_{2}(x)$ implies $h<H_{2}^{-}(x)$ for all $h\in H_{1}(x)$. If $H_{1}(x)\subset H_{2}(x)$ and $H_{2}(x)\subset H_{1}(x)$, then a contradiction is produced, i.e., $H_{1}^{+}(x)<H_{2}^{-}(x)\leqslant H_{2}^{+}(x)<H_{1}^{-}(x)\leqslant H_{1}^{+}(x)$.
@@
(ii) The definition of inclusion relationship in [23] is too specific to have a small range of applications. For example, let $H_{1}(x)=\{0.1,0.2,0.5\}$, $H_{2}(x)=\{0.6,0.7,0.9\}$ and $H_{3}(x)=\{0.1,0.5,0.8\}$, then $H_{1}(x)\subset H_{2}(x)$ is obvious. However, this definition cannot describe the relationship of $H_{2}(x)$ and $H_{3}(x)$ and is not applicable for many cases of hesitant fuzzy sets.
This study introduces various types of inclusion relations for hesitant fuzzy sets determined by the strength of the information contained in hesitant fuzzy membership degrees. Initially, a comprehensive example is provided to clarify Definition 2.10.
@@
Let $U=\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}\}$ be a set of decision-making schemes and $H$ be an expert team that consists of three experts. $H(U)=\frac{\{0.9,0.2\}}{x_{1}}+\frac{\{0.6,0.6,0.5\}}{x_{2}}+\frac{\{0.7,0.5,0.5\}}{x_{3}}+\frac{\{0.8,0.6,0.5\}}{x_{4}}+\frac{\{0.9,0.3,0.1\}}{x_{5}}+\frac{\{0.9,0.8,0.7\}}{x_{6}}$ are the estimated values of schemes provided by experts, in which the estimated values for $x_{1}$ are 0.9 and 0.2 that are obtained through the evaluations made by experts. One of the three experts fails to evaluate the scheme $x_{1}$.
(1) Here, $x_{1}$ has an estimated value of 0.9, which is greater than or equal to all the estimated values for $x_{2}$, it is possible that the scheme $x_{1}$ is better than the scheme $x_{2}$, denoted as $H(x_{2})\subset_{p}H(x_{1})$.
(2) Here, $0.5670.55=mean[H(x_{1})]<mean[H(x_{2})]=$ 0.567, where $mean[\cdot]$ is the mean value operator. In comparing the mean values of estimated values, we find that the scheme $x_{2}$ is better than the scheme $x_{1}$, denoted as $H(x_{1})\subset_{m}H(x_{2})$.
(3) On the one hand, the best estimated value of $x_{3}$ is 0.7, which is greater than or equal to the best estimated value of $x_{2}$. On the other hand, the worst estimated value of $x_{3}$ is 0.5, which is greater than or equal to the worst estimated value of $x_{2}$. To compare the respective best and worst cases of schemes $x_{2}$ and $x_{3}$, it is acceptable that the scheme $x_{3}$ is better than the scheme $x_{2}$, denoted as $H(x_{2})\subset_{a}H(x_{3})$.
(4) To compare the estimated values for schemes $x_{3}$ and $x_{4}$ one by one ($0.7\leqslant 0.8$; $0.5\leqslant 0.6$; and $0.5\leqslant 0.5$), it is strongly credible that the scheme $x_{4}$ is better than the scheme $x_{3}$, denoted as $H(x_{3})\subset_{s}H(x_{4})$.
(5) We can obtain $H(x_{1})\subset_{s}H(x_{5})$ after truncating the tail estimated value of $x_{5}$, i.e., deleting the estimated value 0.1 of $x_{5}$. This case is denoted as $H(x_{1})\subset_{st}H(x_{5})$ and is recorded briefly as $H(x_{1})\subset_{t}H(x_{5})$.
(6) The worst estimated value of $x_{6}$ is greater than or equal to the best estimated value of $x_{3}$. Thus, it is necessary that the scheme $x_{6}$ is better than the scheme $x_{3}$, which is denoted as $H(x_{3})\subset_{n}H(x_{6})$.
Comparing the mean value of estimated values is a common approach for decision-making; however, while doing so, some important information may be lost, such as the best and the worst estimated values.
@@
(2) Let $q$ be a positive integer. To take $q$ larger numbers of $W$ and construct a best $q$-subsequence of $W$, denoted as $W(q)$, i.e., $|W(q)|=q$ and $w\leqslant inf(W(q))$ for $w\in W-W(q)$. For example, let $W=\{0.9,0.8,0.7,0.65,0.6,0.5\}$, when $q=2$, $W(q)=\{0.9,0.8\}$; when $q=3$, $W(q)=\{0.9,0.8,0.7\}$.
@@
(3) Let $q=inf\{|H_{1}(x)|,|H_{2}(x)|\}$, $H_{1}(x)(q)$ and $H_{2}(x)(q)$ be the best $q$-subsequences of $H_{1}(x)$ and $H_{2}(x)$, respectively. Thereafter, $H_{1}(x)(q)\preccurlyeq H_{2}(x)(q)$ if and only if one of $H_{1}(x)\subset_{s}H_{2}(x)$ and $H_{1}(x)\subset_{t}H_{2}(x)$ holds.
(4) Let $K=H_{1}(x)\sqcap H_{2}(x)$, if $sup(H_{1}(x)-K)\leqslant inf(K)\leqslant sup(K)\leqslant inf(H_{2}(x)-K)$, then one of $H_{1}(x)\subset_{s}H_{2}(x)$ and $H_{1}(x)\subset_{t}H_{2}(x)$ holds.
@@
A brief summary for Proposition 2.12 is shown as follows,
{``$\subset_{n}"\Rightarrow``\subset_{s}"\Rightarrow``\subset_{a}"\Rightarrow``\subset_{p}",\\ ``\subset_{n}"\Rightarrow``\subset_{s}"\Rightarrow``\subset_{m}",\\ ``\subset_{n}"\Rightarrow``\subset_{t}"\Rightarrow``\subset_{p}".\\ \end{cases}$
2.3 Some propositions about the intersection and union of two hesitant fuzzy sets
Let $\bar{A}$ and $\bar{B}$ be two classical sets, we know that $\bar{A}\sqcap\bar{B}\sqsubset\bar{A}$, $\bar{A}\sqcap\bar{B}\sqsubset\bar{B}$, $\bar{A}\sqsubset\bar{A}\sqcup\bar{B}$, $\bar{B}\sqsubset\bar{A}\sqcup\bar{B}$, and $\bar{A}\sqcap\bar{B}\sqsubset\bar{A}\sqcup\bar{B}$. However, these rules are not necessarily true for two hesitant fuzzy sets $A$ and $B$ (shown as Propositions 2.13-2.20).